___5___ - ___8___ = 1
2x-3 3-2x

Answer by stanbon(75887)

___5___ - ___8___ = 1
2x-3 3-2x
-------------------------------
Re-write:
[5/(2x-3)] - [8/(3-2x)] = 1
Multiply thru by (2x-3)[-(2x-3)] = -(2x-3)^2
-----------------------------
-5(2x-3) -8(2x-3) = -(2x-3)^2
-10x + 15 -16x + 24 = -[4x^2 -12x + 9]
-26x + 39 = -4x^2 + 12x -9
4x^2 -38x + 48 = 0
2(x^2 - 19x + 24) = 0
x^2 -19x + 24 = 0
x = [19 +- sqrt(19^2-4*1*24)]/2
x = [19 +- sqrt(265)]/2
x = [19 +- 16.2788...]/2
x = 17.639410... or x = 1.3606...
cheers,
Stan H.